# Constraining the Extragalactic Magnetic Field: Auger Data Meet UHECR Propagation Modeling

### v1

@@ Abstract
Recent analyses from the Pierre Auger Collaboration suggest correlations between the arrival directions of Ultra-High-Energy Cosmic Rays (UHECRs) and catalogs of starburst galaxies (SGBs) and jetted active galactic nuclei (AGNs). We revisit these analyses using the same methodology as Auger, but explicitly incorporating UHECR deflections in turbulent extragalactic magnetic fields (EGMFs). We demonstrate that while for SBGs the same sources as for the generic Auger analysis dominate the catalog correlations, jetted AGNs are dominated by Centaurus A when accounting for source distances and deflections. Using our framework, we derive 90% confidence level upper limits on the local EGMF strength of 4.4 nG Mpc1/2 for SBGs and 6.7 nG Mpc1/2 for jetted AGNs. Assuming instead that the UHECR deflections predominantly arise from the Galactic magnetic field (GMF), we obtain a GMF upper limit of $1.4\,\mu$G kpc1/2 for a Galactic halo size of 30 kpc.
@@ 1 Introduction
Cosmic rays are energetic charged ions that arrive at Earth with a range of energies spanning many orders of magnitude up to about $10^{20}$ eV. Theoretical considerations suggest that Ultra-High-Energy Cosmic Rays (UHECRs) are likely produced by extreme energetic events giving rise to high velocity outflows, such as Active Galactic Nuclei (AGNs), Gamma-Ray Bursts (GRBs) or Tidal Disruption Events (TDEs) (Hillas, 1984; Aharonian & Neronov, 2005; Begelman et al., 1984; Farrar & Piran, 2014). Sometimes also Starbust Galaxies (SBGs) are considered (see e.g. Anchordoqui (2018)), although the sub-relativistic velocities of their outflows cast doubt on whether they are able to accelerate particles up to the UHE regime (Romero et al., 2018).
@@
The Pierre Auger Observatory (Auger) (Aab et al., 2015), in the Southern Hemisphere, and the Telescope Array (TA) (Abu-Zayyad et al., 2012), in the Northern Hemisphere, are the two largest detectors of UHECRs. Auger has recently reported the detection of large-scale anisotropy, namely a dipole in the arrival directions of UHECRs with energies above $8$ EeV with a post-trial significance for a departure from isotropy at the level of $6.8\sigma$ (Aab et al., 2017, 2018; Abdul Halim et al., 2024b). The fact that this dipole points away from the Galactic center suggests that the majority of UHECR sources at these energies are of extragalactic, rather than Galactic, origin. Furthermore, the latest joint analysis combining data from both Auger and TA has reached a post-trial significance for a departure from isotropy at the level of $4.6\sigma$ (Abdul Halim et al., 2025b).
@@
At even higher energies, $\gtrsim 40$ EeV, Auger has observed a significant excess of events on smaller angular scales, which appears to correlate with local extragalactic structures. In particular, an excess has been found in either the direction of local SBGs and jetted AGNs (dominated by Cen A), over expectations for an isotropic distribution, at the $4.2\sigma$ and $3.3\sigma$ confidence level, respectively (Aab et al., 2018; Abreu et al., 2022). Even though only a small fraction (about 10%) of the observed UHECRs correlate with the selected objects, the results motivate the study of SBGs and AGNs as UHECR sources.
@@
Uncertainties in the UHECR mass composition, the Galactic magnetic field (GMF), and the extragalactic magnetic field (EGMF) all collectively hinder the identification of the origin of UHECRs. The role of EGMFs is that, in addition to the GMFs, they smear out the location of the UHECR source and wipe out the time correlation, which may be used to identify transients. [^1] For example, for the AGN Cen-A, time delays of Myr are expected for EGMFs with nG strengths (Mbarek & Caprioli, 2025). Understanding EGMFs is, therefore, crucial to understanding the transport of UHECRs.
@@
Although the details about EGMFs remain unclear (Dolag et al., 2005; Vernstrom et al., 2021), we anticipate that measurements of the deflections of the cosmic rays from candidate sources can themselves provide new insights into the EGMF properties. Within local distances from their origin, before becoming isotropic, the UHECR arrival directions retain some angular correlation with their source. This residual angular structure for local sources can provide a valuable probe of the strength of the EGMF, since the extent of deflection, and hence the angular spread around the source position, is directly influenced by the EGMF strength. In addition to magnetic deflections, the density of UHECR sources shapes the observed angular distribution: A lower source density enhances the visibility of local sources, whose anisotropic signatures can be resolved more distinctly, whereas a higher source density leads to a more isotropic background dominated by numerous distant sources. By modeling the interplay between magnetic deflection and source distribution, one can constrain the EGMF more effectively (van Vliet et al., 2021).
@@
Combined fits to both the energy spectrum, composition, and anisotropy can provide additional insights. Such a combined analysis by Auger  (Abdul Halim et al., 2024a), using energy dependent “blurring angles” as a description of the effect of turbulent magnetic field deflections, found preference for a starburst origin of UHECR. Alternatively, Wykes et al. (2018); Taylor et al. (2023) have found that a Cen A origin interpretation of the energy spectrum and small-scale anisotropy is also possible within a scenario in which galaxies possess magnetized giant haloes. A better understanding of the local EGMF is crucial here to discern between these two scenarios.
@@
In this work, we investigate constraints on the local extragalactic magnetic field strength and coherence length on distance scales of the local sources. Our approach closely follows Auger’s work in Abreu et al. (2022) using the same methodology (see Sec. 2.1), but explicitly incorporating the deflections caused by extragalactic magnetic fields, taking source distances into account (see Sec. 2.2). Consequently, in our approach, the signal fraction and magnetic field strength will be treated as key parameters. Our primary goal is to derive constraints for the EGMF strength, see Sec. 3. We will finally discuss our results and conclude in Secs. 4 and 5, respectively.
@@ 2 Methods
This work closely follows the analysis of UHECR data by Auger from Aab. & et al. (2015); Aab et al. (2018); Abreu et al. (2022). Throughout, we assume that the small-scale anisotropy observed by Auger is genuine, using the correlation with candidate sources to constrain the magnetic-field parameters required to produce these UHECR hotspots.
@@ 2.1 Analysis of the Pierre Auger Collaboration
In the papers of Auger, they tested their measured arrival directions for cosmic rays with energies $E>32$ EeV against a blind search for excesses in the sky, autocorrelations, correlations with the Galactic and supergalactic planes, the Galactic center, catalogs of candidate host galaxies, and the Centaurus region. In Abreu et al. (2022), the entire Phase 1 data set of the Pierre Auger Observatory, i.e. all data preceding the upgrade to AugerPrime (Aab et al., 2016), was used for this analysis. That article included a release of the data set itself, as well as the dedicated analysis software and the catalogs of galaxies that were used. The analyses in this work are based on that data set, analysis software, and those catalogs.
@@
For the correlations with catalogs of candidate host galaxies, Auger considered four different catalogs. The catalog of SBGs (based on Lunardini et al. (2019)) includes 44 sources with a radio flux at 1.4 GHz larger than 0.2 Jy from the NRAO VLA Sky Survey (NVSS) (Condon et al., 1998) and Parkes (Calabretta et al., 2014) surveys, and within a distance of $2{-}109$ Mpc. The catalog of jetted AGN has 26 source candidates based on the $\gamma$-ray Fermi-LAT 3FHL AGN catalog (Ajello et al., 2017) within a distance $3{-}241$ Mpc, and with an integral flux between 10 GeV and 1 TeV larger than $3.3\times 10^{-11}\mathrm{cm}^{-2}\mathrm{s}^{-1}$. Besides these two catalogs, Auger also considered a catalog representing the large-scale distribution of matter (2MASS (Skrutskie et al., 2006)) and a catalog of all different types of AGN (based on Oh et al. (2018)). These catalogs were used to create model maps of UHECR arrival directions.
@@
The skymap models generated by Auger (see Fig. 10 of Abreu et al. (2022)) illustrate the expected distribution of the UHECR flux across the sky, based on the cataloged source populations. These models represent the anisotropic distribution of cosmic-ray sources as inferred from the source catalogs. The models are for a cosmic-ray flux originating from the sources listed in the catalogs, with no contribution from a random, isotropic background.
@@
To create the model maps based on the source catalogs, the deflection of cosmic rays in Galactic and extragalactic magnetic fields needs to be taken into account. Auger did this by smearing out the flux distribution of the sources using Fisher distributions centered on the source positions. This is mathematically equivalent to a Gaussian distribution on the sphere (Fisher, 1953). These Fisher distributions are represented by a Fisher smearing angle or search radius $\Theta$. The value of $\Theta$ is the same for all sources in a particular model map, and this value is scanned over to find the best fit to the UHECR data. Effectively, this means that Auger ignored the distance to the sources when estimating the deflection from the sources, implying negligible EGMF deflections and dominant GMF deflections. Our analysis deviates from this approach. Instead of applying Fisher distributions, we simulated the propagation of UHECRs through turbulent EGMF for different magnetic-field strengths. This allows us to obtain an upper limit on the strength of EGMFs when neglecting GMF deflections.
@@ 2.2.1 Extragalactic Magnetic Field Model
We model the EGMF as a purely turbulent field, without any large-scale or structured component. As a result, the deflection direction is isotropic, meaning that a particle with a given deflection angle is equally likely to be deflected in any direction perpendicular to its trajectory. Consequently, a single simulated particle does not correspond to a single arrival direction on the sky, but rather to an entire ring of possible arrival directions, all located at the same angular distance (the deflection angle) from the source.
The turbulent magnetic field in our simulations is modeled with a Kolmogorov power spectrum, with a minimum and maximum turbulence scale of $\lambda_{\min}=120~{}\text{kpc}$ and $\lambda_{\max}=1~{}\text{Mpc}$, respectively, which results in a coherence length of $l_{\mathrm{coh}}=0.256~{}\text{Mpc}$. We conduct our simulations for different magnetic-field strengths, sampled on a logarithmic scale, using the values listed in Table 1.
@@
The characteristics of the magnetic field are parametrized by its root-mean-square strength ($B_{\mathrm{rms}}$) and maximum correlation length ($l_{\mathrm{coh}}$). These are combined into an effective parameter, $\tilde{B}=B_{\mathrm{rms}}\sqrt{l_{\mathrm{coh}}}$, consistent with theoretical expectations for cosmic ray diffusion in turbulent fields (Neronov & Vovk, 2010).
The choice of magnetic-field parameters is consistent with bounds derived from gamma-ray observations, which constrain the root-mean-square field strength to $B_{\mathrm{rms}}\geq 10^{-17}$ G for $l_{\mathrm{coh}}\geq 0.1$ Mpc (see, e.g., Acciari et al. (2023)). The selected range of $\tilde{B}$ spans from scenarios with small deflections to those with significant deflections, for a typical source distance of $\sim 4$ Mpc.
@@ 2.2.2 UHECR Propagation from the Sources
The deflection angles are computed using Monte Carlo simulations with CRPropa 3 (Alves Batista et al., 2022), which propagates UHECRs through the magnetic field model described in Sec. 2.2.1. Importantly, in our setup, the simulated deflection angle depends only on the energy of the particle and the comoving distance to the source, and not on the actual position of the source in the sky.
@@
where $E$ and $Z$ are the energy of the cosmic ray and the particle’s charge at the source, respectively. $R_{\mathrm{max}}$ is the maximum rigidity at the source, $\gamma$ is the spectral index. In this work, we use the best-fit parameters for star-formation rate source evolution obtained from Batista et al. (2019) for $\gamma$ and $R_{\mathrm{max}}$ as: $\gamma=-1.3$, $R_{\mathrm{max}}=10^{18.2}$ V. The source evolution itself will not have a significant effect on our simulations, as we are only simulating the arrival distribution from relatively nearby sources. For the composition, we only include nitrogen nuclei at the source as they typically provide the dominant contribution in the relevant energy range (between 32 EeV and 80 EeV).
In our simulations, we account for all relevant propagation effects for all relevant isotopes, including: pion production, photodisintegration and pair production in ambient low-energy photon fields (the cosmic microwave background (CMB) and the extragalactic background light (EBL) from Gilmore et al. (2012)), as well as adiabatic energy losses and nuclear decay.
@@ 2.3 Model map creation and data comparison
In this analysis, to construct the model map, we generate a ring centered at the original source position with a radius equal to the simulated deflection angle for each simulated particle. We do this by subtracting an inner disk from an outer disk (defining the ring’s thickness), computing its solid angle area, counting the number of HEALPix pixels it covers (using NSIDE = 64, which corresponds to a pixel size on the order of the angular resolution of the Observatory (Górski et al., 2005)), and normalizing each ring based on that number of pixels. All rings associated with a given source are summed and normalized, and the final contribution from each source is reweighted according to its flux in the source catalog. The final probability density map is obtained by summing the weighted contributions from all sources, with the center of the contribution from each source centered at the source position in the sky, effectively capturing the expected distribution of arrival directions on the sky.
In the likelihood analysis, $\alpha$ is varied along with the effective magnetic field strength $\tilde{B}$, which governs the angular spread of events around source positions due to magnetic deflections, to identify the combination of parameters that best fits the observed UHECR arrival directions.
The test statistic is maximized as a function of the signal fraction $\alpha$, in steps of $1\%$, and the effective magnetic field strength $\tilde{B}$, with $l_{\mathrm{coh}}=0.256~{}\text{Mpc}$ and magnetic-field strengths as listed in Table 1. We scan the energy thresholds $E_{\mathrm{th}}$ ranging from 32 to 80 EeV. For each energy threshold, we explore the two-dimensional parameter space defined by the magnetic field strength and the signal fraction.
For a given energy threshold, the test statistic follows a $\chi^{2}$ distribution with two degrees of freedom, as described by Wilks (1938). Consequently, the $1\sigma$ and $2\sigma$ confidence levels (C.L.s) on the best-fit parameters, namely the signal fraction $\alpha$ and the effective magnetic field strength $\tilde{B}$, are determined from iso-TS (constant test statistic) contours in the two-dimensional parameter space. These contours correspond to regions where the TS value is reduced by fixed amounts relative to its maximum. Specifically, the $1\sigma$ and $2\sigma$ confidence regions correspond to TS decreases of 2.3 and 6.2 units, respectively, consistent with the expectations from the $\chi^{2}$ distribution for two parameters (Abreu et al., 2022).

### v2

@@ Abstract
[…]
@@ 1 Introduction
Cosmic rays are energetic charged particles that arrive at Earth with a range of energies spanning many orders of magnitude up to about $10^{20}$ eV. Theoretical considerations suggest that ultra-high-energy cosmic rays (UHECRs) are likely produced by extreme energetic events giving rise to high velocity outflows, such as active galactic nuclei (AGNs), gamma-ray bursts (GRBs) or tidal disruption events (TDEs) (Hillas_1984; Aharonian_2005; RevModPhys.56.255; 2014arXiv1411.0704F). Sometimes also starbust galaxies (SBGs) are considered (see e.g. 2018PhRvD..97f3010A), although the sub-relativistic velocities of their outflows cast doubt on whether they are able to accelerate particles up to the UHE regime (2018A&A...616A..57R).
@@
The Pierre Auger Observatory (Auger) (2015172), in the Southern Hemisphere, and the Telescope Array (TA) (ABUZAYYAD201287), in the Northern Hemisphere, are the two largest detectors of UHECRs. Auger has recently reported the detection of large-scale anisotropy, namely a dipole in the arrival directions of UHECRs with energies above $8$ EeV with a post-trial significance for a departure from isotropy at the level of $6.8\sigma$ (2017; Aab_2018; PierreAuger:2024fgl). The fact that this dipole points away from the Galactic center suggests that the majority of UHECR sources at these energies are of extragalactic, rather than Galactic, origin. Furthermore, the latest joint analysis combining data from both Auger and TA has reached a post-trial significance for a departure from isotropy at the level of $4.6\sigma$ (TelescopeArray:2025yvu).
@@
At even higher energies, $\gtrsim 40$ EeV, Auger has observed a significant excess of events on smaller angular scales, which appears to correlate with local extragalactic structures. In particular, an excess has been found in either the direction of local SBGs and jetted AGNs (dominated by Cen A), over expectations for an isotropic distribution, at the $4.2\sigma$ and $3.3\sigma$ confidence level, respectively (Aab_2018; Abreu_2022). In TelescopeArray:2025yvu, Auger and TA also did a combined analysis of the correlation between UHECR arrival directions and nearby SBGs and found a significance of $4.4\sigma$. Even though only a small fraction (about 10%) of the observed UHECRs correlate with the selected objects, the results motivate the study of SBGs and AGNs as UHECR sources.
@@
Uncertainties in the UHECR mass composition, the Galactic magnetic field (GMF), and the extragalactic magnetic field (EGMF) all collectively hinder the identification of the origin of UHECRs. The role of EGMFs is that, in addition to the GMFs, they smear out the location of the UHECR source and wipe out the time correlation, which may be used to identify transients. [^1] For example, for the AGN Cen-A, time delays of Myr are expected for EGMFs with nG strengths (2025arXiv250201022M). The GMF adds a minimum delay of $\sim$1–50 yr for protons with $E\sim 60$–100 EeV along high-latitude sightlines, and much longer ($\gtrsim 10^{3}$ yr) for heavier nuclei. Even if extragalactic magnetic fields are negligible, this GMF “floor” limits the types of transients that can be temporally associated with UHECRs, with only the highest-energy protons along favorable directions potentially identifiable. Understanding EGMFs is, therefore, crucial to understanding the transport of UHECRs.
@@
Although the details about EGMFs remain unclear (Dolag_2005; Vernstrom2021), we anticipate that measurements of the deflections of the cosmic rays from candidate sources can themselves provide new insights into the EGMF properties. Within local distances from their origin, before becoming isotropic, the UHECR arrival directions retain some angular correlation with their source. This residual angular structure for local sources can provide a valuable probe of the strength of the EGMF, since the extent of deflection, and hence the angular spread around the source position, is directly influenced by the EGMF strength. In addition to magnetic deflections, the density of UHECR sources shapes the observed angular distribution: A lower source density enhances the visibility of local sources, whose anisotropic signatures can be resolved more distinctly, whereas a higher source density leads to a more isotropic background dominated by numerous distant sources. By modeling the interplay between magnetic deflection and source distribution, one can constrain the EGMF more effectively (van_Vliet_2021).
@@
Combined fits to both the energy spectrum, composition, and anisotropy can provide additional insights. Such a combined analysis by Auger  (PierreAuger:2023htc), using energy dependent “blurring angles” as a description of the effect of turbulent magnetic field deflections, found preference for a starburst origin of UHECR. Alternatively, Wykes:2017nno; Taylor:2023qdy have found that a Cen A origin interpretation of the energy spectrum and small-scale anisotropy is also possible within a scenario in which galaxies possess magnetized giant haloes. A better understanding of the local EGMF is crucial here to discern between these two scenarios.
@@
In this work, we investigate constraints on the local extragalactic magnetic field strength and coherence length on distance scales of the local sources. Our approach closely follows Auger’s work in Abreu_2022 using the same methodology (see Sec. 2.1), but explicitly incorporating the deflections caused by extragalactic magnetic fields, taking source distances into account (see Sec. 2.2). Consequently, in our approach, the signal fraction and magnetic field strength will be treated as key parameters. Our primary goal is to derive constraints for the EGMF strength, see Section. 3. We will finally discuss our results and conclude in Secs. 4 and 5, respectively.
@@ 2 Methods
This work closely follows the analysis of UHECR data by Auger from Aab_2015; Aab_2018; Abreu_2022. Throughout, we assume that the small-scale anisotropy observed by Auger is genuine, using the correlation with candidate sources to constrain the magnetic-field parameters required to produce these UHECR hotspots.
@@ 2.1 Analysis of the Pierre Auger Collaboration
In the papers of Auger, they tested their measured arrival directions for cosmic rays with energies $E>32$ EeV against a blind search for excesses in the sky, autocorrelations, correlations with the Galactic and supergalactic planes, the Galactic center, catalogs of candidate host galaxies, and the Centaurus region. In (Abreu_2022), the entire Phase 1 data set of the Pierre Auger Observatory, i.e. all data preceding the upgrade to AugerPrime (PierreAuger:2016qzd), was used for this analysis. That article included a release of the data set itself, as well as the dedicated analysis software and the catalogs of galaxies that were used. The analyses in this work are based on that data set, analysis software, and those catalogs.
@@
For the correlations with catalogs of candidate host galaxies, Auger considered four different catalogs. The catalog of SBGs (based on Lunardini_2019) includes 44 sources with a radio flux at 1.4 GHz larger than 0.2 Jy from the NRAO VLA Sky Survey (NVSS) (Condon_1998A) and Parkes (Calabretta2014) surveys, and within a distance of 21092{-}109 Mpc. The catalog of jetted AGN has 26 source candidates based on the $\gamma$-ray Fermi-LAT 3FHL AGN catalog (Fermi-LAT_2017), and with an integral flux between 10 GeV and 1 TeV larger than $3.3\times 10^{-11}\mathrm{cm}^{-2}\mathrm{s}^{-1}$ within a distance 32413{-}241 Mpc. Besides these two catalogs, Auger also considered a catalog representing the large-scale distribution of matter (2MASS (2MASS:2006qir)) and a catalog of all different types of AGN (based on Oh:2018wzc). These catalogs were used to create model maps of UHECR arrival directions.
@@
The skymap models generated by Auger (see Fig. 10 of Abreu_2022) illustrate the expected distribution of the UHECR flux across the sky, based on the cataloged source populations. These models represent the anisotropic distribution of cosmic-ray sources as inferred from the source catalogs. The models are for a cosmic-ray flux originating from the sources listed in the catalogs, with no contribution from a random, isotropic background.
@@
To create the model maps based on the source catalogs, the deflection of cosmic rays in Galactic and extragalactic magnetic fields needs to be taken into account. Auger did this by smearing out the flux distribution of the sources using Fisher distributions centered on the source positions. This is mathematically equivalent to a Gaussian distribution on the sphere (Fisher1953). These Fisher distributions are represented by a Fisher smearing angle or search radius $\Theta$. The value of $\Theta$ is the same for all sources in a particular model map, and this value is scanned over to find the best fit to the UHECR data. Effectively, this means that Auger ignored the distance to the sources when estimating the deflection from the sources, implying negligible EGMF deflections and dominant GMF deflections. Our analysis deviates from this approach. Instead of applying Fisher distributions, we simulated the propagation of UHECRs through turbulent EGMF for different magnetic-field strengths. This allows us to obtain an upper limit on the strength of EGMFs when neglecting GMF deflections.
@@ 2.2 Modeling cosmic ray deflections
We model the EGMF as a purely turbulent field, without any large-scale or structured component. The turbulent magnetic field in our simulations is modeled with a Kolmogorov power spectrum, with a minimum and maximum turbulence scale of $\lambda_{\min}=120~\text{kpc}$ and $\lambda_{\max}=1~\text{Mpc}$, respectively, resulting in a coherence length of $l_{\mathrm{coh}}=0.256~\text{Mpc}$. $\lambda_{\max}$ is chosen so that the resulting coherence length is significantly smaller than the typical distance to nearby sources, ensuring that UHECRs experience multiple independent deflections (Batista_2019). $\lambda_{\min}$ is chosen so that the lowest energy particles considered still have resonant modes to interact with, even for the strongest magnetic fields. In the simulations, the magnetic-field modes are assigned random directions, while the magnitude of the field strength in each mode follows a power law as described in 1999ApJ...520..204G. We conduct our simulations for different magnetic-field strengths, sampled on a logarithmic scale, using the values listed in Table 1.222Table 1 gives $B_{\rm rms}$ rather than $\tilde{B}$ because $B_{\rm rms}$ is the direct input parameter used in the CRPropa simulations. The magnetic-field strengths were sampled on a logarithmic scale, with a finer step size in the middle of the parameter range to capture the transition region. While $B_{\rm rms}$ provides a clear, model-independent reference, $\tilde{B}$ is the effective quantity that directly impacts UHECR propagation and is discussed in the results.
@@
The characteristics of the magnetic field are parametrized by its root-mean-square strength ($B_{\mathrm{rms}}$) and maximum correlation length ($l_{\mathrm{coh}}$). These are combined into an effective parameter, $\tilde{B}=B_{\mathrm{rms}}\sqrt{l_{\mathrm{coh}}}$, consistent with theoretical expectations for cosmic ray diffusion in turbulent fields (Neronov_2010).
The choice of magnetic-field parameters is consistent with bounds derived from gamma-ray observations, which constrain the root-mean-square field strength to $B_{\mathrm{rms}}\geq 10^{-17}$ G for $l_{\mathrm{coh}}\geq 0.1$ Mpc (see e.g. MAGIC2023). The selected range of $\tilde{B}$ spans from scenarios with small deflections to those with significant deflections, for a typical source distance of $\sim 4$ Mpc.
@@ 2.2.2 UHECR propagation from the sources
The deflection angles are computed using Monte Carlo simulations with CRPropa 3 (AlvesBatista_2022), which propagates UHECRs through the magnetic field model described in Section. 2.2.1. Importantly, in our setup, the simulated deflection angle depends only on the energy of the particle and the comoving distance to the source, and not on the actual position of the source in the sky.
We simulated $5\times 10^{5}$ particles per source using CRPropa, propagating them through our turbulent extragalactic magnetic field setup. The source sends out particles isotropically. We track all particles out to a sphere centered at the source, with a radius equal to the source distance. We evaluate the deflection angles through the comparison of the emission and observer momentum vectors. All particles arriving at the observer sphere are used to create the image for one specific source. In this way, we effectively average over the turbulent magnetic field structure orientation. Our approach ignores any coherent shifts produced by the turbulent field realization in a particular direction (Dolgikh:2025bac). This “large observer” approach mitigates the extremely low detection probability of a point-like observer while preserving the directional information relevant for subsequent analysis.
@@
where $E$ and $Z$ are the energy of the cosmic ray and its charge at the source, respectively. $R_{\mathrm{max}}$ is the maximum rigidity at the source, $\gamma$ is the spectral index. In this work, we use the best-fit parameters for star-formation rate source evolution obtained from Batista_2019 for $\gamma$ and $R_{\mathrm{max}}$ as: $\gamma=-1.3$, $R_{\mathrm{max}}=10^{18.2}$ V. The source evolution itself will not have a significant effect on our simulations, as we are only simulating the arrival distribution from relatively nearby sources. For the composition, we only include nitrogen nuclei at the source as they typically provide the dominant contribution in the relevant energy range (between 32 EeV and 80 EeV), see Section. 4 for further discussion and justification of these assumptions.
In our simulations, we account for all relevant propagation effects for all relevant isotopes, including: pion production, photodisintegration and pair production in ambient low-energy photon fields (the cosmic microwave background (CMB) and the extragalactic background light (EBL) from Gilmore_2012), as well as adiabatic energy losses and nuclear decay.
@@ 2.3 Model map creation and data comparison
In this analysis, to construct the image map for each single particle, we rotate its simulated deflection angle around its original direction vector. This generates a ring centered at the original source position with a radius equal to the simulated deflection angle. In this way, we gain in efficiency as each simulated particle generates a ring instead of a single position in the sky. A source image is created by summing the rings, each normalized to its corresponding surface, of all simulated particles. The resultant image is approximately Gaussian around the source position. The combination of all source images, normalized according to their fluxes in the source catalog, gives the full simulated catalogue sky map. This is created on a HEALPix map using NSIDE = 64, which corresponds to a pixel size on the order of the angular resolution of the Observatory (Górski_2005).
In the likelihood analysis, the signal fraction $\alpha$ is varied along with the effective magnetic field strength $\tilde{B}$, which governs the angular spread of events around source positions due to magnetic deflections, to identify the combination of parameters that best fits the observed UHECR arrival directions.
The test statistic is maximized as a function of $\alpha$, in steps of $1\%$, and the effective magnetic field strength $\tilde{B}$, with $l_{\mathrm{coh}}=0.256~\text{Mpc}$ and magnetic-field strengths as listed in Table 1. We scan the energy thresholds $E_{\mathrm{th}}$ ranging from 32 to 80 EeV. For each energy threshold, we explore the two-dimensional parameter space defined by the magnetic field strength and the signal fraction.
For a given energy threshold, the test statistic follows a $\chi^{2}$ distribution with two degrees of freedom, as described by Wilks. Consequently, the $1\sigma$ and $2\sigma$ confidence levels (C.L.s) on the best-fit parameters, namely the signal fraction $\alpha$ and the effective magnetic field strength $\tilde{B}$, are determined from iso-TS (constant test statistic) contours in the two-dimensional parameter space. These contours correspond to regions where the TS value is reduced by fixed amounts relative to its maximum.
Specifically, the $1\sigma$ and $2\sigma$ confidence regions correspond to TS decreases of 2.3 and 6.2, respectively, reflecting joint confidence regions for two parameters rather than the one-parameter intervals (TS decreases of 1 and 4) typically used for marginalized or profiled distributions, consistent with the expectations from the $\chi^{2}$ distribution for two parameters (Abreu_2022).